answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What does the standard error mean?

For a sample of data it is a measure of the spread of the observations about their mean value.

What is probability of a shot child?

If you mean the odds of a child in the US being shot- actually, very low- and has been getting lower for several years. Some groups with a political agenda claim there are more deaths from children being shot than there are ALL deaths of children- so check your sources carefully- and see how they define "Child". Assuming child means a human under 18 years old, there about 74.5 million of them in the US (some groups consider a 26 year old to be a child) The FBI Uniform Crime Report shows 454 murders of children, with about 20% being shot. That is about 1 in 818,000. However, the odds of a child being shot goes up if you live in a high crime city, belong to a gang, or commit criminal acts.

Experimental probability of rolling the sum of 2?

That means that you should roll a die many times, count how often you get the number "2", then divide this by the total number of rolls.

If the die is "fair" (no extra weight on one side), you would expect this experimental probability to be somewhere close to the theoretical probability of 1/6, at least, if you roll often enough.

If a card is drawn at random from an ordinary deck of cards what is the probability that it is a diamond or a four?

number of cards in a deck =52

number of cards that are diamond =13

number of cards that are four =4

number of cards that are diamond and four =1

Probability that the card drawn is a diamond or a four=

13/52+4/52 -1/52

=16/52 or 4/13

How do you get your class size in probability and statistics?

One determinant is what the data will be used for, whether they need to be combined with other data for further analyses. If, for example, you want to study human mortality, you may want to use the same class sizes as for population statistics - which are usually 5-year age bands. That will enable you to calculate deaths rates. However, you may additionally wish to have a band for "less than 1 year" and "1 to 5 years" to allow you to study infant mortality in greater detail. At the other end, you may wish to group together "90 and above".

If there are no such constraints, you should divide up the range of values into 6-10 classes. Ideally, no class should have very few (<5%) observations. If there are, then combine adjoining classes - there is no requirement for the class size to be the same across the whole range.

One determinant is what the data will be used for, whether they need to be combined with other data for further analyses. If, for example, you want to study human mortality, you may want to use the same class sizes as for population statistics - which are usually 5-year age bands. That will enable you to calculate deaths rates. However, you may additionally wish to have a band for "less than 1 year" and "1 to 5 years" to allow you to study infant mortality in greater detail. At the other end, you may wish to group together "90 and above".

If there are no such constraints, you should divide up the range of values into 6-10 classes. Ideally, no class should have very few (<5%) observations. If there are, then combine adjoining classes - there is no requirement for the class size to be the same across the whole range.

One determinant is what the data will be used for, whether they need to be combined with other data for further analyses. If, for example, you want to study human mortality, you may want to use the same class sizes as for population statistics - which are usually 5-year age bands. That will enable you to calculate deaths rates. However, you may additionally wish to have a band for "less than 1 year" and "1 to 5 years" to allow you to study infant mortality in greater detail. At the other end, you may wish to group together "90 and above".

If there are no such constraints, you should divide up the range of values into 6-10 classes. Ideally, no class should have very few (<5%) observations. If there are, then combine adjoining classes - there is no requirement for the class size to be the same across the whole range.

One determinant is what the data will be used for, whether they need to be combined with other data for further analyses. If, for example, you want to study human mortality, you may want to use the same class sizes as for population statistics - which are usually 5-year age bands. That will enable you to calculate deaths rates. However, you may additionally wish to have a band for "less than 1 year" and "1 to 5 years" to allow you to study infant mortality in greater detail. At the other end, you may wish to group together "90 and above".

If there are no such constraints, you should divide up the range of values into 6-10 classes. Ideally, no class should have very few (<5%) observations. If there are, then combine adjoining classes - there is no requirement for the class size to be the same across the whole range.

What is the significance of probable cause?

"Probable Cause" is the foundatiion on which all arrests are based. It is a reasonable belief that a person has committed a crime. The test for whether probable cause exists for the purpose of an arrest is whether facts and circumstances within the officer's knowledge are sufficient to warrant a prudent person to believe a suspect has committed, is committing, or is about to commit a crime. See: http://www.lectlaw.com/def2/p089.htm

How many cards are in a yugiho deck?

I do 60 but I'm not 100% sure if it's right

40-60 cards is the limit.

Odds of rolling 6 ones with 10 dice?

The probability of rolling 6 ones with 10 dice is: 10C6 (1/6)6 (5/6)4 = 2.170635034...

x 10-3 ≈ 0.22%

What are the Odds of a woman having a baby on her birthday?

1 out of 365

These odds improve if she has sex and gets pregnant about 3 months after her birthday.

WHAT IS PROBABILITY OF A DECK OF CARDS DRAWING A TWO OF CLUBS AND JACK OF HEARTS?

The answer depends on

  • how many cards are drawn,
  • whether of not they are drawn at random, and
  • whether or not the cards are replaced before the next card is drawn.
There is no information on these and so it is not possible to give a definitive answer to the question.

How many 3 cards are in a deck?

There are four 3's in a standard deck of 52 cards.

What is the probability of rolling a 2 on the first dice and an odd number on the second dice?

None, because you cannot have the first or second dice: it is the first die or second die.

The probability is 1/6 * 1/2 = 1/12

Which method is least appropriate for performing a simulation to estimate probability?

I believe that stimulation is always inappropriate. Simulation, maybe, but that was not what the question was about.

What artist did a painting of three men in red coats and powered wigs playing cards in front of a fire?

I believe it was Frank Moss Bennett. I have a reproduction of this painting you are talking about. When I look through several on-line galleries I have not seen this particular one under his name. His signature is on the bottom left side.

What does Chupacabra mean?

Chupacabra is the Spanish word for "goatsucker" because it is a creature that sucks the blood out of animals.

It means that it sucks ("chupa") the blood from goats ("cabras"). Other small animal deaths have been attributed to the chupacabras. Although there is no scientific support for its existence, the belief in the chupacabra is strong in some areas, due to its extensive folklore.

What is the probability that a specified seeding has no neighbors within 1 meter?

The answer will depend on the density of the seedlings and their distribution - in 2-D space..

Trending Questions
Why is it important that the scientific community use a standard set of units in measuring and recording data? How many prime numbers in a deck of cards? Can someone have two heads? What is the probability of rolling a four on a number cube? How many ways can the letters tjkl be arranged? There are 52 playing cards they are divided into 13 suits diamonds hearts clubs and spades What is the probability of picking random hearts or clubs what is the answer in fraction decimal? Why can't you use FOIL to multiply a binomial? A model that helps explain the results of Mendel's crosses is called a? What is the probability of NOT landing on a 2 of a circle that is numbered from 1 to 8? What is the area under normal distribution curve between z1.50 and z2.50? When two probabilities are multiplied the probability represents a compound event true or false? What is the probability that the offspring of a homozygous dominant individual and a homozygous recessive individual exhibit the dominant phenotype? The chance that a giving will occur usually expressed between the number 0 will not and 1 the event it will occur? What is the difference between natural infinite and real infinite set? Are the diagonals of a trapezium perpendicular? What is the standard deviation of a set data in which all the data values are the same? Is binomial distribution is characterized by situations that are analogous to coin tossing? Which is greater -0.3 or -0.001? What is the probability of rolling a seven with two dice given that at least one of them is a four? How many hearts and diamonds are there in a deck of playing cards?